xx ReductionTest

ReductionTest

Christian D. Ott <cott@aei.mpg.de>

April 18 2006

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11domain10Scope: shared from GRIDx01KEYWORD

11

7 Interfaces

General Implements: reductiontest2

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  redvars  10  x01  compact110
11                 10  redvar  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels113
11                 10  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/ReductionTest2. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:11x
11redvars[3]11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x com2xlocal x setup integrands for com calculation x

11x11After: 11whiskyxinitial
11 x 11Language:11c
11 x 11Sync:  11redvars
11 x 11Type: 11function
11 x 11Writes: 11reductiontest2::redvar(everywhere)

11

CCTKxEVOL x com2xlocal x setup integrands for com calculation x

11x11After: 11molxevolution
11 x 11Language:11c
11 x 11Sync:  11redvars
11 x 11Type: 11function

11

CCTKxPOSTREGRIDINITIAL x com2xlocal x setup integrands for com calculation x

11x11Language:11c
11 x 11Sync:  11redvars
11 x 11Type: 11function

11

CCTKxPOSTREGRID x com2xlocal x setup integrands for com calculation

11x11Language:11c
11 x 11Sync:  11redvars
11 x 11Type: 11function